# Vortices and Factorization

### v3

@@ 15 Conclusions and Open Problems
$\tau_{n}\Delta\tau_{n+1}-2(\nabla\tau_{n+1}\cdot\nabla\tau_{n})+\tau_{n+1}\Delta\tau_{n}=0,\quad\tau_{0}=1,$ (123)
where $\Delta$ and $\nabla$ denote two-dimensional Laplacian and gradient, respectively. This chain is called harmonic [10], because at the first step $\Delta\tau_{1}=0$, i.e., $\tau_{1}$ is a harmonic function, and $\psi_{n}=\tau_{n+1}/\tau_{n}$ are natural generalizations of harmonic functions [^34] For example, $\psi_{n}$ satisfy quadrature identities in “harmonic” quadrature domains (see e.g. [33, 34]).. The harmonic chain can be generalized to any number of dimensions: Non-terminating solutions of (123) are $\tau$-functions for potentials of algebraically integrable Schrodinger operators (see e.g. [10], [17]). One can view (123) as a generalization of the Darboux chain (35), that starts from the free Schrodinger operator [^35] A chain that starts from $u_{0}\not=0$ has the form $\tau_{n}\Delta\tau_{n+1}-2(\nabla\tau_{n+1}\cdot\nabla\tau_{n})+\tau_{n+1}\Delta\tau_{n}-u_{0}\tau_{n}\tau_{n+1}=0$, $\tau_{0}=1$., to any number of dimensions. In one dimension, the Adler-Moser polynomials constitute complete set of non-terminating solutions of the chain. In two dimensions, the soliton related solutions exhaust all non-terminating solutions in the class of homogeneous polynomials: The family of two-dimensional solutions consists of an infinite number of branches labelled by $0<k_{1}<k_{2}<\dots$, rather than a single sequence, as in one dimension. Complete classification of non-terminating solutions in all dimensions is a hard open problem (see e.g. [10, 17]) [^36] It is conjectured that families listed in [17] constitute complete solution of the problem..
16 Appendix 1: Main Sequences and Polynomial $\tau$-functions: Wronskian and Pfaffian Representations.
@@
$s_{1}=-3t_{3},\,s_{2}=45t_{5},\,s_{3}=-1575t_{7},\,s_{4}=99255\left(t_{9}-\frac{t_{3}}{3}\right),\,s_{5}=9823275(t_{3}^{2}t_{5}-t_{11}),\,\dots$
@@
$L_{2}=\partial_{z}^{3}-6\,{\frac{\left(5\,{z}^{8}+80\,{z}^{3}r_{{1}}-10\,{z}^{4}s_{{2}}+s_{2}^{2}\right)}{\left({z}^{5}-4\,r_{{1}}+s_{{2}}z\right)^{2}}}\partial_{z},$
@@
Note that the transformation $L_{0}\to L_{n}$ can be performed through a chain of permutations and re-factorizations, involving the first-order operators $\partial_{z}-V_{i}$, where $V_{i}$ are rational (see e.g. [14, 26]). As we saw in section 9, in contrast to the $\Lambda=1$ case, intertwining identities cannot, in general, be constructed solely by permuting factors in $L_{n}$: The factorization chain now includes intermediate iterations. For instance, for $n>0$, the order of the intertwining operator $T_{n}$ is $23n-$ 2, i.e., ${\rm ord}(T_{n+1})-{\rm ord}(T_{n})=3$. This is because each transformation $L_{n}\to L_{n+1}$, $n>0$, presented by (81,84), actually corresponds to three steps, i.e., to three permutations and re-factorizations in the chain ($L_{0}\to L_{1}$ is performed in a single step). As a consequence, the operators ${\cal T}_{n}$ that intertwine $L_{n}$ and $L_{n+1}$:
@@
are of the third order, when $n>0$ (of the first order when $n=0$).

### v4

@@ Abstract
$\tau_{n}\Delta\tau_{n+1}-2(\nabla\tau_{n+1}\cdot\nabla\tau_{n})+\tau_{n+1}\Delta\tau_{n}=0,\quad\tau_{0}=1,\quad\nabla=(\partial_{X},\partial_{Y}),$ (123)
where $\Delta=\nabla\cdot\nabla$ and $\nabla$ denote two-dimensional Laplacian and gradient, respectively. This chain is called harmonic [10], because at the first step $\Delta\tau_{1}=0$, i.e., $\tau_{1}$ is a harmonic function, and Schrodinger operators corresponding to $\tau_{n}$ are natural generalizations of the Laplace operator [33, 34]. The harmonic chain can be generalized to any number of dimensions: Non-terminating solutions of (123) are $\tau$-functions for potentials of algebraically integrable Schrodinger operators (see e.g. [10], [17]). One can view (123) as a generalization of the Darboux chain (36), that starts from the free Schrodinger operator [^34] A chain that starts from $u_{0}\not=0$ has the form $\tau_{n}\Delta\tau_{n+1}-2(\nabla\tau_{n+1}\cdot\nabla\tau_{n})+\tau_{n+1}\Delta\tau_{n}-u_{0}\tau_{n}\tau_{n+1}=0$, $\tau_{0}=1$., to any number of dimensions. In one dimension, the Adler-Moser polynomials constitute complete set of non-terminating solutions of the chain. In two dimensions, the soliton related solutions exhaust all non-terminating solutions in the class of homogeneous polynomials: The family of two-dimensional solutions consists of an infinite number of branches labelled by $0<k_{1}<k_{2}<\dots$, rather than a single sequence, as in one dimension. Complete classification of non-terminating solutions in all dimensions is a hard open problem (see e.g. [10, 17]) [^35] It is conjectured that families listed in [17] constitute complete solution of the problem..
Appendix A Main Sequences and Polynomial $\tau$-functions: Wronskian and Pfaffian Representations.
@@
$s_{1}=-3t_{3},\,s_{2}=45t_{5},\,s_{3}=-1575t_{7},\,s_{4}=99255\left(t_{9}-\frac{t_{3}}{3}\right),\,s_{5}=9823275\left(\frac{4}{3}t_{3}^{2}t_{5}-t_{11}\right),\,\dots$
@@
$L_{2}=\partial_{z}^{3}-6\,{\frac{5\,{z}^{8}+80\,{z}^{3}r_{{1}}-10\,{z}^{4}s_{{2}}+s_{2}^{2}}{\left({z}^{5}-4\,r_{{1}}+s_{{2}}z\right)^{2}}}\partial_{z},$
@@
Note that the transformation $L_{0}\to L_{n}$ can be performed through a chain of permutations and re-factorizations, involving the first-order operators $\partial_{z}-V_{i}$, where $V_{i}$ are rational (see e.g. [14, 26]). As we saw in section 9, in contrast to the $\Lambda=1$ case, intertwining identities cannot, in general, be constructed solely by permuting factors in $L_{n}$: The factorization chain now includes intermediate iterations. For instance, for $n>0$, the order of the intertwining operator $T_{n}$ is $23n-$ 2, i.e., ${\rm ord}(T_{n+1})-{\rm ord}(T_{n})=3$. This is because each transformation $L_{n}\to L_{n+1}$, $n>0$, presented by (81,84), actually corresponds to three steps, i.e., to three permutations and re-factorizations in the chain ($L_{0}\to L_{1}$ is performed in a single step):
$L_{n}=B_{n}A_{n}\to A_{n}B_{n}=\tilde{B}_{n}\tilde{A}_{n}\to\tilde{A}_{n}\tilde{B}_{n}=\accentset{\approx}{B}_{n}\accentset{\approx}{A}_{n}\to\accentset{\approx}{A}_{n}\accentset{\approx}{B}_{n}=B_{n+1}A_{n+1}=L_{n+1},\quad n>0.$
As a consequence, the operators ${\cal T}_{n}=\accentset{\approx}{A}_{n}\tilde{A}_{n}A_{n}$, $n>0$ that intertwine $L_{n}$ and $L_{n+1}$:
@@
are of the third order, when $n>0$ (of the first order when $n=0$, ${\cal T}_{0}=A_{0}$). For example
${\cal T}_{1}=\partial_{z}^{3}-{\frac{8\left(r_{{1}}+{z}^{5}\right)}{\left({z}^{5}-4\,r_{{1}}+s_{{2}}z\right)z}}\partial_{z}^{2}+2\,{\frac{9\,{z}^{5}-3\,s_{{2}}z+4\,r_{{1}}}{\left({z}^{5}-4\,r_{{1}}+s_{{2}}z\right){z}^{2}}}\partial_{z},\quad{\cal T}_{1}L_{1}=L_{2}{\cal T}_{1}.$